# Dielectric environment engineering via 2D material heterostructure formation on hybrid photonic crystal nanocavity

### v1

@@ Abstract
Hybrid integration of two-dimensional (2D) materials with nanophotonic platforms has enabled compact optoelectronic devices by leveraging the unique optical and electronic properties of atomically thin layers. While most efforts have focused on coupling 2D materials to pre-defined photonic structures, the broader potential of 2D heterostructures for actively engineering the photonic environment remains largely unexplored. In our previous work, we employed single types of 2D material and showed that even monolayer flakes can locally induce high- nanocavities in photonic crystal (PhC) waveguides through effective refractive index modulation. Here, we extend this concept by demonstrating that further transferring of 2D material flakes onto the induced hybrid nanocavity to form heterostructures enable more flexibility for post-fabrication dielectric environment engineering of the cavity. We show that the high- hybrid nanocavities remain robust under sequential flake stacking. Coupling optically active MoTe2 flake to these cavities yields enhanced photoluminescence and reduced emission lifetimes, consistent with Purcell-enhanced light-matter interactions. Additionally, encapsulation with a top hBN layer leads to a significant increase in the cavity factor, in agreement with numerical simulations. Our results show that these heterostructure stacks not only preserve the cavity quality but also introduce an additional degrees of control — via flake thickness, refractive indices, size and interface design — offering a richer dielectric environment modulation landscape than what is achievable with monolayers alone, providing a versatile method toward scalable and reconfigurable hybrid nanophotonic systems.
@@ I Introduction
Despite being mechanically compatible with the target substrates, the 2D material integration process inevitably perturbs the surrounding optical environment due to the change in the local dielectric environment caused by the material, potentially introducing undesirable effects such as shifting of optical resonance frequency, modifying mode profile and emission properties, scattering and absorption optical losses. Nonetheless, the influence of 2D materials on the local dielectric environment is usually simply ignored or regarded as minor perturbation to be compensated by adjusting the parameter of the photonic structure. Yet, the ability of 2D materials to modify their dielectric surroundings presents an underexplored opportunity. Rather than being viewed as a detriment, the integration of 2D materials onto photonic structures can be harnessed to engineer the dielectric environment, facilitating additional means for light–matter coupling.
@@ II Design and numerical simulation of hybrid nanocavity
To guide our choice of flake materials and thicknesses to form the hybrid nanocavity and subsequent multi-layer stacking studies, we first perform finite-difference time-domain (FDTD) simulations [45] of single-flake hybrid nanocavity to determine the factor as a function of flake thickness for three 2D materials of different refractive indices: hBN (), WSe2 () and MoTe2 (), using dielectric constants obtained from Ref. [46]. In these simulations, the flake width is kept at and is assumed to cover the PhC structure completely along the -direction. The PhC waveguide consists of 24 (14) air holes along the – (–) direction, with and . The factor exhibits a peak at a specific flake thickness—approximately 3 nm for MoTe2, 5 nm for WSe2 and 20 nm for hBN. These trends suggest the existence of an optimal condition for light confinement, determined by both the refractive index and thickness of the flake (Fig. 1c).
@@
In-plane losses, on the other hand, arise mainly from scattering and leakage at the lateral edges of the flake and from field escaping through the ends of the PhC waveguide. Since the cavity is induced by the flake, the flake width essentially governs the lateral extend fo the mode field profile. The simulated mode intensity profiles of the fundamental, i.e. lowest energy, mode of a a single-flake hybrid nanocavity — considering a hBN flake (marked yellow) with width of 14 and thickness of 10 nm — is shown in Fig. 1d. This broad spatial distribution brings the field closer to the waveguide ends, facilitating in-plane leakage, degrading optical confinement. A combination of thin and wide flake would therefore result in poor out-of-plane and in-plane confinement causing low cavity factor, corresponding to the leftmost region of Fig. 2c.
@@
We then consider the case of a double-flake hybrid nanocavity, with the simulated fundamental mode intensity profile shown in Fig. 1d. On top of the hBN flake (width of 14 and thickness of 10 nm), we consider another flake (cyan) of MoTe2 with width of 10 and thickness of 5 nm. The centers of the flakes and PhC are aligned. The mode profile remains largely unchanged aside from a decrease in its lateral extend since the second flake has a smaller width than the first. For the given parameters, the factor decreases from about for the single-hBN-flake case to when the addition of the MoTe2 flake. Since we do not consider any absorption in the FDTD simulations, the drop in factor can simply be attributed to the change in the refractive index profile. We also investigated various combinations and configurations of 2D material flakes—for example, using WSe2 and hBN as the first and second flakes, respectively, or varying the relative flake sizes (Supplementary Fig. S3 and S4). We find that cavity formation and optical confinement are primarily governed by the in-plane and out-of-plane losses arising from the refractive index profile determined by the specific flake stacking configuration, as well as the PhC waveguide parameters.
@@ III Fabrication and optical spectroscopy
To demonstrate the utility of such hybrid nanocavity for light-matter coupling with additional stacking of 2D materials, a multilayer MoTe2 flake is subsequently transferred onto the hBN-hybrid nanocavity (Fig. 2d). The width of the flake along the waveguide is about 16. The PL emission from the MoTe2/hBN-hybrid nanocavity under excitation power of of 0.01 mW is shown in Fig. 2e. Despite the much lower excitation power and the indirect bandgap of the multilayer MoTe2 flake, it still exhibits relatively bright exciton emission, with the PL spectrum consisting of practically no emission from silicon and only from MoTe2. The broad emission spanning from 1100 nm to 1400 nm allows us to deduce that the MoTe2 flake is of 5 layers thick. The cavity peak around 1310 nm is still clearly visible in the spectrum. From the laser transmission spectrum (Fig. 2f), the cavity peak is found to have broadened, corresponding to . Based on FDTD simulation results presented in Fig. 1d which does not consider absorption losses, the observed reduction in the experimentally measured can be largely attributed to the change in the dielectric environment, and partially by the absorption losses introduced by the MoTe2 flake.
@@
After hBN encapsulation, the hBN/MoTe2/hBN heterostructure stack is now composed mostly of hBN, with an overall thickness of about 35 nm — which is close to 25 nm optimal thickness for high factor as described in a previous paragraph — and therefore leading to the observed increase of the Q-factor. Intuitively, the hBN encapsulation layer acts as a high-quality dielectric spacer, mitigating abrupt refractive index variations and promoting a more homogeneous dielectric environment around the cavity, reducing scattering losses while preserving cavity factor.
@@
Although the fabricated hybrid nanocavities contain 96 air holes along the – direction, our experimental observations align more closely with simulation results of a smaller nanocavities with 24 air holes along the – direction. To investigate this discrepancy, we performed additional finite element method (FEM) simulations of larger PhC waveguides — with 48 (14) air holes along the – (–) direction, which is as large as our computational resources allow — and with flake dimensions matching those of the fabricated devices. Simulation results reveal that an optimal hBN thickness of 5 nm yields the highest factor in this case (Supplementary Fig. S2). Consequently, the simulated factor show a decreasing trend upon the addition of a second MoTe2 flake and subsequent hBN encapsulation since any the overall thicknesses are beyond the optimal value (Supplementary Fig. S4). These results suggest that the effective lateral size of the fabricated hybrid nanocavity is smaller than its actual extent, likely due to in-plane losses caused by scattering or leakage along the PhC waveguide due to fabrication imperfections. This in turn implies that, aside from the waveguide geometry, disorder management offers an additional degree of freedom for engineering the in-plane optical losses, and thus tuning the optimal flake thickness required to achieve high-quality optical confinement.
@@
We further performed time-resolved PL measurements to to probe the exciton recombination dynamics in MoTe2, both in its bare form (away from the PhC region) and when coupled to the hybrid cavity, with and without hBN encapsulation. For these measurements, a pulsed Ti:sapphire laser is used for excitation at 780 nm. The PL emission collected by the objective lens is then filtered spectrally with a long-pass filter before being coupled into an optical fiber connected to a superconducting nanowire single photon detector. A comparison of the decay curves of the PL from the cavity and bare MoTe2 in the MoTe2/hBN hybrid nanocavity before and after hBN-encapsulation is summarized in Fig. 3. The decay curves exhibit a fast component associated with radiative recombination of bright excitons and trions [47]. This fast component is of particular interest, as the cavity Purcell effect selectively enhances the radiative emission of bright excitons by increasing the local photonic density of states. In contrast, slow components in the decay curves are typically attributed to thermally activated recombination of dark excitons [47], which are not significantly influenced by the cavity. Therefore, we will focus our discussion on the fast decay time.
@@
In contrast, the hBN-encapsulated MoTe2/hBN hybrid nanocavity sample shows overall shorter lifetimes with the bare MoTe2 emission giving t1 = ps. The decay times of the cavity emission is reduced further to t1 = ps. The lifetime reduction of both the emission from the bare MoTe2 and the cavity rules out any cavity effect. Previous observation of such lifetime reduction in WSe2 emission lead to the suggestion that hBN-encapsulation reduces both the radiative and non-radiative lifetimes equally [48]. We rule out any linewidth reduction due to hBN encapsulation [49, 50, 51] (which would lead to the counter-effect of increased lifetime) since there is significant phonon broadening at room temperature at which our measurements are performed. We also do not expect any lifetime reduction facilitated by microcavity formation with hBN encapsulation at such elevated temperatures [52]. It is possible that MoTe2 has degraded over time, leading to increasing defects and thus causing an overall decrease in the lifetime throughout the flake. The observed reduction in the lifetimes could also be due to hBN encapsulation which suppresses doping and environmental charge fluctuations, leading to an increased proportion of neutral excitons with shorter lifetimes, over trions (charged excitons) which tend to have longer lifetimes [50, 53, 47, 52, 51]. The extracted lifetimes gives Purcell enhancement factor of 3.8.

### v2

@@ Abstract
Hybrid integration of two-dimensional (2D) materials with nanophotonic platforms has enabled compact optoelectronic devices by leveraging the unique optical and electronic properties of atomically thin layers. While most efforts have focused on coupling 2D materials to pre-defined photonic structures, the broader potential of 2D heterostructures for actively engineering the photonic environment remains largely unexplored. In our previous work, we employed single types of 2D material and showed that even monolayer flakes can locally induce high- nanocavities in photonic crystal waveguides through effective refractive index modulation. Here, we extend this concept by demonstrating that further transferring of 2D material flakes onto the induced hybrid nanocavity to form heterostructures enable more flexibility for post-fabrication dielectric environment engineering of the cavity. We show that the high- hybrid nanocavities remain robust under sequential flake stacking. Coupling optically active MoTe2 flake to these cavities yields enhanced photoluminescence and reduced emission lifetimes, consistent with Purcell-enhanced light-matter interactions. Additionally, encapsulation with a top hBN layer leads to a significant increase in the cavity factor, in agreement with numerical simulations. Our results show that these heterostructure stacks not only preserve the cavity quality but also introduce an additional degrees of control — via flake thickness, refractive indices, size and interface design — offering a richer dielectric environment modulation landscape than what is achievable with monolayers alone, providing a versatile method toward scalable and reconfigurable hybrid nanophotonic systems.
@@ I Introduction
Despite being mechanically compatible with the target substrates, the 2D material integration process inevitably perturbs the optical environment due to changes in the local dielectric profile caused by the 2D material itself. This could introduce undesirable effects such as the shifting of optical resonance frequency, modifying mode profile and emission properties, as well as increased optical losses. Nonetheless, the influence of 2D materials on the local dielectric environment is usually simply ignored or regarded as minor perturbation to be compensated by adjusting the parameter of the photonic structure. Yet, the ability of 2D materials to modify their dielectric surroundings presents an underexplored opportunity. Rather than being viewed as a detriment, the integration of 2D materials onto photonic structures can be harnessed to engineer the dielectric environment, facilitating additional means for light–matter coupling.
@@ II Design and numerical simulation of hybrid nanocavity
To guide our choice of flake materials and thicknesses to form the hybrid nanocavity and subsequent multi-layer stacking studies, we first perform finite-difference time-domain (FDTD) simulations [45] of single-flake hybrid nanocavity to determine the factor as a function of flake thickness for three 2D materials of different refractive indices: hBN (), WSe2 () and MoTe2 (), using dielectric constants obtained from Ref. 46. In these simulations, the flake width is kept at and is assumed to cover the PhC structure completely along the -direction. The PhC waveguide consists of 24 (14) air holes along the – (–) direction, with and . The factor exhibits a peak at a specific flake thickness—approximately 3 nm for MoTe2, 5 nm for WSe2 and 20 nm for hBN. These trends suggest the existence of an optimal condition for light confinement, determined by both the refractive index and thickness of the flake (Fig. 1c).
@@
In-plane losses, on the other hand, arise mainly from scattering and leakage at the lateral edges of the flake and from field escaping through the ends of the PhC waveguide. Since the cavity is induced by the flake, the flake width essentially governs the lateral extent fo the mode field profile. The simulated mode intensity profiles of the fundamental, i.e. lowest energy, mode of a a single-flake hybrid nanocavity — considering a hBN flake (marked yellow) with width of 14 and thickness of 10 nm — is shown in Fig. 1d. This broad spatial distribution brings the field closer to the waveguide ends, facilitating in-plane leakage, degrading optical confinement. A combination of thin and wide flake would therefore result in poor out-of-plane and in-plane confinement causing low cavity factor, corresponding to the leftmost region of Fig. 2c.
@@
We then consider the case of a double-flake hybrid nanocavity, with the simulated fundamental mode intensity profile shown in Fig. 1d. On top of the hBN flake (width of 14 and thickness of 10 nm), we consider another flake (cyan) of MoTe2 with width of 10 and thickness of 5 nm. The centers of the flakes and PhC are aligned. The mode profile remains largely unchanged aside from a decrease in its lateral extend since the second flake has a smaller width than the first. For the given parameters, the factor decreases from about for the single-hBN-flake case to with the addition of the MoTe2 flake. Since we do not consider any absorption in the FDTD simulations, the drop in factor can simply be attributed to the change in the refractive index profile. We also investigated various combinations and configurations of 2D material flakes—for example, using WSe2 and hBN as the first and second flakes, respectively, or varying the relative flake sizes (Supplementary Fig. S3 and S4). We find that cavity formation and optical confinement are primarily governed by the in-plane and out-of-plane losses arising from the refractive index profile determined by the specific flake stacking configuration, as well as the PhC waveguide parameters.
@@ III Fabrication and optical spectroscopy
To demonstrate the utility of such hybrid nanocavity for light-matter coupling with additional stacking of 2D materials, a multilayer MoTe2 flake is subsequently transferred onto the hBN-hybrid nanocavity (Fig. 2d). The width of the flake along the waveguide is about 16. The PL emission from the MoTe2/hBN-hybrid nanocavity under excitation power of 0.01 mW is shown in Fig. 2e. Despite the much lower excitation power and the indirect bandgap of the multilayer MoTe2 flake, it still exhibits relatively bright exciton emission, with the PL spectrum consisting of practically no emission from silicon and only of emission from MoTe2. The broad emission spanning from 1100 nm to 1400 nm allows us to deduce that the MoTe2 flake is of 5 layers thick. The cavity peak around 1310 nm is still clearly visible in the spectrum. From the laser transmission spectrum (Fig. 2f), the cavity peak is found to have broadened, corresponding to . Based on FDTD simulation results presented in Fig. 1d which does not consider absorption losses, the observed reduction in the experimentally measured can be largely attributed to the change in the dielectric environment, and partially by the absorption losses introduced by the MoTe2 flake.
@@
After hBN encapsulation, the hBN/MoTe2/hBN heterostructure stack is now composed mostly of hBN, with an overall thickness of about 35 nm — which is close to 25 nm optimal thickness for high factor as described in a previous paragraph — and therefore leading to the observed increase of the -factor. Intuitively, the hBN encapsulation layer acts as a high-quality dielectric spacer, mitigating abrupt refractive index variations and promoting a more homogeneous dielectric environment around the cavity, reducing scattering losses while preserving cavity factor.
@@
Although the fabricated hybrid nanocavities contain 96 air holes along the – direction, our experimental observations align more closely with simulation results of a smaller nanocavities with 24 air holes along the – direction. To investigate this discrepancy, we performed additional finite element method (FEM) simulations of larger PhC waveguides — with 48 (14) air holes along the – (–) direction, which is as large as our computational resources allow — and with flake dimensions matching those of the fabricated devices. Simulation results reveal that an optimal hBN thickness of 5 nm yields the highest factor in this case (Supplementary Fig. S2). Consequently, the simulated factor shows a decreasing trend upon the addition of a second MoTe2 flake and subsequent hBN encapsulation since any the overall thicknesses are beyond the optimal value (Supplementary Fig. S4). These results suggest that the effective lateral size of the fabricated hybrid nanocavity is smaller than its actual extent, likely due to in-plane losses caused by scattering or leakage along the PhC waveguide due to fabrication imperfections. This in turn implies that, aside from the waveguide geometry, disorder management offers an additional degree of freedom for engineering the in-plane optical losses, and thus tuning the optimal flake thickness required to achieve high-quality optical confinement.
@@
We further performed time-resolved PL measurements to probe the exciton recombination dynamics in MoTe2, both in its bare form (away from the PhC region) and when coupled to the hybrid cavity, with and without hBN encapsulation. For these measurements, a pulsed Ti:sapphire laser is used for excitation at 780 nm. The PL emission collected by the objective lens is then filtered spectrally with a long-pass filter before being coupled into an optical fiber connected to a superconducting nanowire single photon detector. A comparison of the decay curves of the PL from the cavity and bare MoTe2 in the MoTe2/hBN hybrid nanocavity before and after hBN-encapsulation is summarized in Fig. 3. The decay curves exhibit a fast component associated with radiative recombination of bright excitons and trions [47]. This fast component is of particular interest, as the cavity Purcell effect selectively enhances the radiative emission of bright excitons by increasing the local photonic density of states. In contrast, slow components in the decay curves are typically attributed to thermally activated recombination of dark excitons [47], which are not significantly influenced by the cavity. Therefore, we will focus our discussion on the fast decay time.
@@
In contrast, the hBN-encapsulated MoTe2/hBN hybrid nanocavity sample shows overall shorter lifetimes with the bare MoTe2 emission giving t1 = ps. The decay times of the cavity emission is reduced further to t1 = ps. The lifetime reduction of both the emission from the bare MoTe2 and the cavity rules out any cavity effect. Previous observation of such lifetime reduction in WSe2 emission lead to the suggestion that hBN-encapsulation reduces both the radiative and non-radiative lifetimes equally [48]. We rule out any linewidth reduction due to hBN encapsulation [49, 50, 51] (which would lead to the counter-effect of increased lifetime) since there is significant phonon broadening at room temperature at which our measurements are performed. We also do not expect any lifetime reduction facilitated by microcavity formation with hBN encapsulation at such elevated temperatures [52]. It is possible that MoTe2 has degraded over time, leading to increasing defects and thus causing an overall decrease in the lifetime throughout the flake. The observed reduction in the lifetimes could also be due to hBN encapsulation which suppresses doping and environmental charge fluctuations, leading to an increased proportion of neutral excitons with shorter lifetime, over trions (charged excitons) which tend to have longer lifetimes [50, 53, 47, 52, 51]. The extracted lifetimes give Purcell enhancement factor of 3.8.